A 0.299 g sample of a gas was held in a 0.220 L container at 25.0°C and its pressure
was determined to be 0.757 atm. What is the molar mass of the gas?
1
Expert's answer
2020-12-06T10:33:37-0500

PV = nRT


PV = "\frac{w}{m}\times RT"


M = "\frac{wRT}{Pv}"


M= "\frac{0.299 \times 0.0821 \times 298}{0.757\times 0.22}"


M= 43.92 g/mol